$5400 is invested, part of it at 12% and part of it at 8%. For a certain year, the total yield is $ 552.00. How much was invested at each rate?
Answer by TimothyLamb(4379)

x = amount invested at 12%
y = amount invested at 8%
---
x + y = 5400
0.12x + 0.08y = 552
---
put the system of linear equations into standard form
---
x + y = 5400
0.12x + 0.08y = 552
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = amount invested at 12% = 3000
y = amount invested at 8% = 2400
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---